# Global Continuation of Stable Periodic Orbits in Systems of Competing Predators

### v1

@@ Abstract
We develop a continuation technique to obtain global families of stable periodic orbits, delimited by transcritical bifurcations at both ends. We formulate a zero-finding problem whose zeros correspond to families of periodic orbits. We then define a Newton-like fixed-point operator and establish its contraction near a numerically computed approximation of the family. To verify the contraction, we derive sufficient conditions expressed as inequalities on the norms of the fixed-point operator, and involving the numerical approximation. These inequalities are then rigorously checked by the computer via interval arithmetic. To show the efficacy of our approach, we prove the existence of global families in an ecosystem with Holling’s type II functional response, and thereby solve a stable connection problem proposed by Butler and Waltler in 1981. Our method does not rely on restricting the choice of parameters and is applicable to many other systems that numerically exhibit global families.
@@ 1 Introduction
The purpose of this article is to present an effective continuation method using computer-assisted proofs to obtain global families of periodic orbits, undergoing transcritical bifurcations at invariant boundary planes. Our interest in this question originated from studying the following system (see [28, 29]) that describes the population dynamics of two predators $X_{j}(t)$ for $j=1,2$ and one prey $S(t)$:
@@
with initial conditions in +$\mathbb{R}^{3}_{+}:=\{(X_{1},X_{2},S)\,:\,X_{1}>0,\,X_{2}>0,\,S>0\}$. The system (1) involves ten positive parameters: $\kappa$ is the carrying capacity of the prey, $\gamma$ is its intrinsic rate of increase, and, for the $j$-th predator, $m_{j}$ is its maximum birth rate, $a_{j}$ is its half-saturation constant, $d_{j}$ is its death rate, and $y_{j}$ is its yield conversion factor. The functional response in (1) is called Holling’s type II, which is also known as the Michaelis–Menten kinetics in chemistry; see [10, 34]. For experimental results on (1), see [25] and the references therein.
Since $\mathbb{R}_{+}^{3}$ is invariant under the dynamics of the system (1), solutions with initial conditions in $\mathbb{R}_{+}^{3}$ remain positive, that is, $(X_{1}(t),X_{2}(t),S(t))\in\mathbb{R}^{3}_{+}$ for all $t\in\mathbb{R}$. Biodiversity described by (1) occurs when all species coexist such that $\liminf_{t\to\infty}X_{j}(t)>0$ for $j=1,2$ and $\liminf_{t\to\infty}S(t)>0$. In other words, coexistence is characterized by the survival of both predators and their prey, thereby offering an alternative perspective to the competitive exclusion principle; see [44].
Numerical results have long suggested that coexistence in +$\mathbb{R}^{3}_{+}$ may manifest through periodic orbits or more intricate dynamics. However, only a few results have been analytically proven, and the existence of periodic orbits is based on strong assumptions about the parameters. We now comment on the relevant literature. First, the dynamics near both invariant boundary planes
$Q_{1}:=\{X_{1},0,S)\,:\,X_{1}>0,\,S>0\},\qquad Q_{2}:=\{(0,X_{2},S)\,:\,X_{2}>0,\,S>0\}$ (2)
@@
$E_{1}:=\left(\frac{\gamma y_{1}(\lambda_{1}+a_{1})(\kappa-\lambda_{1})}{\kappa m_{1}},0,\lambda_{1}\right),\qquad\text{ where }\qquad\lambda_{1}:=\frac{1}{\kappa}\left(\frac{a_{1}d_{1}}{m_{1}-d_{1}}\right),$ (3)
as we assume $\kappa-\lambda_{1}>0$ and $m_{1}-d_{1}>0$. Moreover, $E_{1}$ undergoes a local Hopf bifurcation in $Q_{1}$ when $02\lambda_{1}+a_{1}-\kappa=$ 0, triggering a boundary limit cycle $\mathcal{C}_{1}\subset Q_{1}$ for $02\lambda_{1}+a_{1}-\kappa<$ 0; see [47]. Then, positive periodic orbits can bifurcate from $\mathcal{C}_{1}$ via a local transcritical bifurcation; see [9, 10, 47]. Notice that such periodic solutions are established only in a neighborhood of $Q_{1}$ in +$\mathbb{R}^{3}_{+}$, and therefore for a small population size of $X_{1}$. Since (1) remains unchanged by interchanging the index $j=1,2$, we can define $E_{2}$, $\lambda_{2}$, and $\mathcal{C}_{2}$ analogously. Second, geometric singular perturbation theory is applicable for sufficiently large $\gamma\gg 0$ resulting in a positive periodic orbit; see [41]. Third, perturbing a conserved quantity by considering two small difference assumptions on the parameters, $0<a_{2}-a_{1}\ll 1$ and $0<\lambda_{2}-\lambda_{1}\ll 1$, yield stable positive periodic orbits far from both boundary planes $Q_{1}$ and $Q_{2}$; see [34]. Last, with only one small difference assumption, $0<\lambda_{2}-\lambda_{1}\ll 1$, a local hybrid Hopf bifurcation occurs by eliminating a line of equilibria, also yielding stable positive periodic orbits far from both boundary planes; see [43]. We emphasize that all existing results are either inherently local or rely on restricting the choice of parameters.
In contrast, in this article, we present a continuation method to prove a family of positive periodic orbits, which is global in the sense that it connects two boundary limit cycles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. The family corresponds to a curve in the ten-dimensional parameter space. Following [29], we parameterize such a curve by the carrying capacity $\kappa>0$. Moreover, the family consists of stable periodic orbits in +$\mathbb{R}^{3}_{+}$, meaning that the associated Floquet exponents of each periodic orbit are 00, $\mu_{1}$, and $\mu_{2}$ such that $\mathrm{Re}(\mu_{1})<0$ and $\mathrm{Re}(\mu_{2})<0$. Our continuation method does not rely on restricting the choice of parameters, due to the nature of the computer-assisted proof developed in Sections 23. As an application, we choose the following set of parameter values (noticing that $y_{1},y_{2},\gamma$ can always be rescaled to $1$)
@@
    $\mathcal{P}(\hat{\kappa}_{1})=\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the boundary limit cycle in $Q_{1}:=\{(X_{1},0,S)\,:\,X_{1}>0,\,S>0\}$;
@@
    $\mathcal{P}(\hat{\kappa}_{2})=\mathcal{C}_{2}$, where $\mathcal{C}_{2}$ is the boundary limit cycle in $Q_{2}:=\{(0,X_{2},S)\,:\,X_{2}>0,\,S>0\}$.
@@
Theorem 1.1 is an affirmative answer to the stable connection problem proposed by Butler and Waltman [10, page 309]:
@@
Numerical evidence (see [29] for instance) hints at a broad parameter region that supports stable positive periodic orbits. However, aside from largely perturbative or asymptotic results, there has not been much progress in solving the problem.
@@
We prove Theorem 1.1 using a state-of-the-art continuation method (see [6, 7]), which the author Hénot recently applied to resolve Marchal’s conjecture in the three-body problem [11]. The continuation of periodic orbits is formalized as a zero-finding problem. To continue through the local transcritical bifurcations from the boundary limit cycles, we build an auxiliary system to (1) whose periodic orbits coincide. Periodic solutions are expressed as Fourier series, where the dependency on the continuation parameter $\kappa>0$ is expressed by expanding the Fourier coefficients as an infinite series of Chebyshev polynomials of the first kind. Thus, a family of periodic orbits corresponds to a zero of an appropriate map $F$ defined on a Banach space of rapidly decaying Fourier–Chebyshev coefficients. Then, the existence of the zero, thereby of the family of periodic orbits, is proved through a contraction argument. The proof reduces to verifying inequalities involving norm estimates, expressed in terms of a numerical approximation of the branch of periodic orbits. The computer is employed to perform a finite, albeit large, number of calculations together with interval arithmetic to prevent rounding errors. The proof of stability is done after the existence of the family, and is based on a similar strategy.
The past two decades have seen a growing interest in computer-assisted proofs (abbr. CAPs) to study dynamical systems. A notable result is the chaotic attractor in the Lorenz system [42]. The computer was involved in both the discovery of chaotic dynamics (non-rigorous computations), and eventually the proof of existence of the chaotic attractor [48] (rigorous computations). CAPs based on a posteriori validation have a long history, going back to pioneering works in the mid-1980s on the Feigenbaum conjectures [21, 22, 40]. Our method is inspired by [51], the infinite-dimensional Krawczyk operator [23], and the approach proposed in [18, 31, 46]. Note that functional analytic methods of CAPs for studying periodic solutions to differential equations date back to the work of Cesari on Galerkin projections [13, 14]. The interested readers are referred to the survey papers [24, 36, 45, 49].
The proof of Theorem 1.1 uses several new techniques in CAPs, notably:
@@
    We use a blow-up (as in “zoom in”) approach to isolate zeros; see also [11, 16, 50]. This allows us to desingularize the continuation as the family undergoes a transcritical bifurcation on the boundary planes $Q_{1}$ and $Q_{2}$.
@@
    We employ the Chebyshev continuation procedure (see [6, 7]) to prove the global family of periodic orbits, and auxiliary data (periods and amplitudes), as a single isolated zero of a map. This is in contrast to the more classical approach of using several uniform contractions; see [5].
@@
    Our proof is efficient in the sense that we handle non-polynomial nonlinearities directly (in the vein of [8]), rather than using a so-called polynomial embedding [26].
@@
This article is structured as follows. In Section 2, we present the rigorous continuation method, which proves the existence part of Theorem 1.1. In Section 3, we address the stability of periodic orbits, where some details are deferred to Appendix A. Finally, in Section 4, we outline potential future work.

### v2

@@ Abstract
We develop a continuation technique to obtain global families of stable periodic orbits, delimited by transcritical bifurcations at both ends. We formulate a zero-finding problem whose zeros correspond to families of periodic orbits. We then define a Newton-like fixed-point operator and establish its contraction near a numerically computed approximation of the family. To verify the contraction, we derive sufficient conditions expressed as inequalities on the norms of the fixed-point operator, and involving the numerical approximation. These inequalities are then rigorously checked by the computer via interval arithmetic. To show the efficacy of our approach, we prove the existence of global families in an ecosystem with Holling’s type II functional response, and thereby solve a stable connection problem proposed by Butler and Waltman in 1981. Our method does not rely on restricting the choice of parameters and is applicable to many other systems that numerically exhibit global families.
@@ 1 Introduction
Understanding stable coexistence in ecological systems remains a central challenge in mathematical biology. To address this, we present an effective continuation method, based on computer-assisted proofs, to obtain global families of stable periodic orbits undergoing transcritical bifurcations at invariant boundary planes. This framework provides a general tool to study the competitive exclusion principle and to rigorously explore mechanisms of stable coexistence; see [28, 56].
More specifically, our interest in this question originated from studying the following system (see [36, 35]) that describes the population dynamics of two predators $X_{j}(t)$ for $j=1,2$ and one prey $S(t)$:
@@
with initial conditions in $\mathbb{R}^{3}_{+}\coloneqq\{(X_{1},X_{2},S)\,:\,X_{1}>0,\,X_{2}>0,\,S>0\}$. The system (1) involves ten positive parameters: $\kappa$ is the carrying capacity of the prey, $\gamma$ is its intrinsic rate of increase, and, for the $j$-th predator, $m_{j}$ is its maximum birth rate, $a_{j}$ is its half-saturation constant, $d_{j}$ is its death rate, and $y_{j}$ is its yield conversion factor. The functional response in (1) is called Holling’s type II, which is also known as the Michaelis–Menten kinetics in chemistry; see [15, 42]. For experimental results on (1), see [30] and the references therein.
Since $\mathbb{R}_{+}^{3}$ is invariant under the dynamics of the system (1), solutions with initial conditions in $\mathbb{R}_{+}^{3}$ remain positive, that is, $(X_{1}(t),X_{2}(t),S(t))\in\mathbb{R}^{3}_{+}$ for all $t\in\mathbb{R}$. Biodiversity described by (1) occurs when all species coexist such that $\liminf_{t\to\infty}X_{j}(t)>0$ for $j=1,2$ and $\liminf_{t\to\infty}S(t)>0$. In other words, coexistence is characterized by the survival of both predators and their prey, thereby disproving the competitive exclusion principle for the system (1).
Numerical results have long suggested that coexistence in $\mathbb{R}^{3}_{+}$ may manifest through periodic orbits or more intricate dynamics. However, only a few results have been analytically proven, and the existence of periodic orbits is based on strong assumptions about the parameters. We now comment on the relevant literature. First, the dynamics near both invariant boundary planes
$Q_{1}\coloneqq\{(X_{1},0,S)\,:\,X_{1}>0,\,S>0\},\qquad Q_{2}\coloneqq\{(0,X_{2},S)\,:\,X_{2}>0,\,S>0\}$ (2)
@@
$E_{1}\coloneqq\left(\frac{\gamma y_{1}(\kappa\lambda_{1}+a_{1})(1-\lambda_{1})}{m_{1}},0,\kappa\lambda_{1}\right),\qquad\text{ where }\qquad\lambda_{1}\coloneqq\frac{1}{\kappa}\left(\frac{a_{1}d_{1}}{m_{1}-d_{1}}\right),$ (3)
as we assume $1-\lambda_{1}>0$ and $m_{1}-d_{1}>0$. Moreover, $E_{1}$ undergoes a local Hopf bifurcation in $Q_{1}$ when $2\kappa\lambda_{1}+a_{1}-\kappa=0$, triggering a boundary limit cycle $\mathcal{C}_{1}\subset Q_{1}$ for $2\kappa\lambda_{1}+a_{1}-\kappa<0$; see [60]. Then, positive periodic orbits can bifurcate from $\mathcal{C}_{1}$ via a local transcritical bifurcation; see [14, 15, 60]. Notice that such periodic solutions are established only in a neighborhood of $Q_{1}$ in $\mathbb{R}^{3}_{+}$, and therefore for a small population size of $X_{1}$. Since (1) remains unchanged by interchanging the index $j=1,2$, we can define $E_{2}$, $\lambda_{2}$, and $\mathcal{C}_{2}$ analogously. Second, geometric singular perturbation theory is applicable for sufficiently large $\gamma\gg 0$ resulting in a positive periodic orbit; see [52]. Third, perturbing a conserved quantity by considering two small difference assumptions on the parameters, $0<(a_{2}-a_{1})/\kappa\ll 1$ and $0<\lambda_{2}-\lambda_{1}\ll 1$, yield stable positive periodic orbits far from both boundary planes $Q_{1}$ and $Q_{2}$; see [42]. Last, with only one small difference assumption, $0<\lambda_{2}-\lambda_{1}\ll 1$, a local hybrid Hopf bifurcation occurs by eliminating a line of equilibria, also yielding stable positive periodic orbits far from both boundary planes; see [53]. We emphasize that all existing results are either inherently local or rely on restricting the choice of parameters.
In contrast, in this article, we present a continuation method to prove a family of positive periodic orbits, which is global in the sense that it connects two boundary limit cycles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. The family corresponds to a curve in the ten-dimensional parameter space. Following [35], we parameterize such a curve by the carrying capacity $\kappa>0$. Moreover, the family consists of stable periodic orbits in $\mathbb{R}^{3}_{+}$, meaning that the associated Floquet exponents of each periodic orbit are 0, $\mu_{1}$, and $\mu_{2}$ such that $\mathrm{Re}(\mu_{1})<0$ and $\mathrm{Re}(\mu_{2})<0$. Our continuation method does not rely on restricting the choice of parameters, due to the nature of the computer-assisted proof developed in Sections 23. As an application, we choose the following set of parameter values (noticing that $y_{1},y_{2},\gamma$ can always be rescaled to $1$)
@@
    $\mathcal{P}(\hat{\kappa}_{1})=\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the boundary limit cycle in $Q_{1}\coloneqq\{(X_{1},0,S)\,:\,X_{1}>0,\,S>0\}$;
@@
    $\mathcal{P}(\hat{\kappa}_{2})=\mathcal{C}_{2}$, where $\mathcal{C}_{2}$ is the boundary limit cycle in $Q_{2}\coloneqq\{(0,X_{2},S)\,:\,X_{2}>0,\,S>0\}$.
@@
Theorem 1.1 is an affirmative answer to the stable connection problem proposed by Butler and Waltman [15, page 309]:
@@
Numerical evidence (see [35] for instance) hints at a broad parameter region that supports stable positive periodic orbits. However, aside from largely perturbative or asymptotic results, there has not been much progress in solving the problem, except for [42], in which the results rely on restricting the small difference assumptions on the parameters $0<(a_{2}-a_{1})/\kappa\ll 1$ and $0<\lambda_{2}-\lambda_{1}\ll 1$.
@@
We prove Theorem 1.1 using a rigorous continuation method, recasted as a single zero-finding problem $F(\chi)=0$. This map $F$ incorporates an auxiliary system to (1) whose periodic orbits coincide, allowing us to continue through the local transcritical bifurcations from the boundary limit cycles. We employ a spectral method, so that periodic solutions are expressed as Fourier series, where the dependency on the continuation parameter $\kappa>0$ is expressed by expanding the Fourier coefficients as an infinite series of Chebyshev polynomials of the first kind. Then, the existence of the zero of $F$, thereby of the family of periodic orbits, is proved by a local contraction argument. The set of sufficient conditions for the appropriate fixed-point operator to be a local contraction amounts to satisfying inequalities relating norm estimates, expressed in terms of a numerical approximation of the branch of periodic orbits. The computer is employed to perform large number needed to verify the contraction hypotheses of calculations together with interval arithmetic to prevent rounding errors. While the strategy detailed here is aimed at addressing the connection problem, an other application was also made in celestial mechanics by one of the author of this article to resolve Marchal’s conjecture in the three-body problem [16]. The proof of stability is done after the existence of the family, and uses a similar contraction strategy, it also utilizes the analytic regularity of the solution curve inferred from the functional framework.
Computer-assisted proofs (abbr. CAPs) based on a posteriori validation go back to the mid-1980s with the works on the Feigenbaum conjectures [25, 26, 48]. The method followed in this article is inspired by a long history of techniques, such as [18, 19, 22, 27, 41, 58, 67]. A notable result of application in the field of dynamical systems, is the proof of the chaotic attractor in the Lorenz system [54]; the computer was involved in both the discovery and, ultimately, the proof of its existence [63]. For a more thorough account on the development of CAPs in the recent years, the interested readers are referred to the survey papers [29, 44, 57, 65]. See also [55] for an introductory account on proving the existence and continuation of periodic orbits using CAPs.
The proof of Theorem 1.1 uses several new techniques in CAPs (see also Table 1), notably:
@@
    We use a blow-up approach to isolate zeros; see also [16, 20, 64]. This allows us to desingularize the continuation as the family undergoes a transcritical bifurcation on the boundary planes $Q_{1}$ and $Q_{2}$.
@@
    We employ a Chebyshev continuation procedure (see [9]) to prove the global family of periodic orbits, and auxiliary data (periods and amplitudes), as a single isolated zero of a map. See [4, 10] for related works using high-order interpolation of solution curves. This is in contrast to the more classical approach of using a piecewise-linear approximation of the solution curve, requiring several uniform contractions; see, e.g., [3, 8, 11, 66].
@@
    Our proof is efficient in the sense that we handle non-polynomial nonlinearities directly (in the vein of [12]), rather than using a so-called polynomial embedding [33], which results in studying vector fields in higher dimensions.
Table 1: Non-exhaustive comparison of CAP strategies.
Theme Alternative CAP strategies Present approach
Bifurcations Blow-up for Hopf bifurcations [20, 64] Blow-up for transcritical bifurcations of periodic orbits from invariant subsets
Global continuation Piecewise-linear approximation of family needing multiple uniform contractions [3, 8, 11, 66], or Taylor expansion of the family [4] Chebyshev expansion of the family
Nonlinearities Polynomial embedding [33] Direct treatment
@@
This article is structured as follows. In Section 1.1, we present a general class of ecological systems to investigate the competitive exclusion principle and highlight the effectiveness of our CAP method. In Section 1.2, we provide a general framework to desgingularize transcritical bifurcations. In Section 2, we present the rigorous continuation method, which proves the existence part of Theorem 1.1. In Section 3, we address the stability of periodic orbits, where some details are deferred to Appendix A. Finally, in Section 4, we outline potential future work.
1.1 Further Applications: the competitive exclusion principle
A central paradigm in ecology is the competitive exclusion principle [28, 56], which states that different species competing for the exact same limited resource cannot stably coexist. In other words, if their ecological niches are identical, one species will eventually exclude the others. While this principle provides a useful heuristic in ecology, there is also considerable mathematical interest in rigorously establishing conditions under which it holds or fails. Notably, H.L. Smith’s work on monotone dynamical systems [61] characterizes settings where competitive interactions enforce exclusion, and where cooperative or more complex structures may allow coexistence. In more applied domains, explicit examples where the principle fails are considered significant contributions in their own right (see e.g., [59]), which demonstrates failure in the context of certain autocatalytic networks. One natural way to demonstrate the failure of the competitive exclusion principle is by establishing the existence of positive periodic orbits, which guarantee ongoing coexistence. In this context, our CAP framework provides a robust method to establish global continuation of such periodic orbits, beyond perturbation arguments. The framework itself opens a principled path toward identifying and rigorously verifying instances of coexistence.
To exemplify, we present the problem of resource competition in Lotka–Volterra systems; see the pioneering work [49]. Suppose that there are $M$ competing predator species consuming $N$ prey resources. The problem requires to determine all pairs $(M,N)\in\mathbb{N}^{2}$ with $M\geq N$ such that the competitive exclusion principle fails, thus supporting biodiversity. Specifically, let $X_{j}(t)$ for $j=1,\dots,M$ and $S_{k}(t)$ for $k=1,\dots,N$ denote the population density of the $j$-th predator and $k$-th prey, respectively. We consider the following general system that describes the population dynamics:
$\begin{cases}\displaystyle\dot{X}_{j}=\left(g_{j}(S_{1},\dots,S_{N})-d_{j}\right)X_{j},&j=1,\dots,M,\\ \displaystyle\dot{S}_{k}\hskip 5.69046pt=h_{k}(S_{k})-\sum_{\ell=1}^{M}c_{k\ell}\,g_{\ell}(S_{1},\dots,S_{N})\,X_{\ell},&k=1,\dots,N.\end{cases}$ (5)
Here, for the $j$-th predator, $g_{j}:\mathbb{R}^{N}_{+}\to\mathbb{R}_{+}$ is the conversion rate to its biomass and $d_{j}>0$ is its death rate. For the $k$-th prey, $h_{k}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is the growth rate when the predation is absent, and $c_{k\ell}\in(0,1)$, which satisfies $\sum_{k=1}^{N}c_{kj}=1$ for each $j=1,\dots,M$, is the fraction of the $\ell$-th predator consuming it.
To choose $g_{j}(S_{1},\dots,S_{N})$ and $h_{k}(S_{k})$, the literature [5, 38, 40, 50, 51] assumes the so-called law of the minimum
$g_{j}(S_{1},\dots,S_{N})=\min_{i=1,\dots,N}\{f_{ij}(S_{i})\},$ (6)
where $f_{ij}(S_{i})$ is of Holling’s type II. Note that the system (1) corresponds to the special case $(M,N)=(2,1)$. The problem of resource competition is uncharted for the logistic quadratic growth
$h_{k}(S_{k})=\gamma(S_{k}^{0}-S_{k})S_{k},$ (7)
where $\gamma>0$ is the intrinsic rate of increase and $S_{k}^{0}>0$ is the growth limiting constant. For the chemostat affine growth
$h_{k}(S_{k})=\gamma(S_{k}^{0}-S_{k}),$ (8)
partial solutions are available. For the cases $(M,N)=(2,2),(3,2)$, the competitive exclusion principle holds; see [50]. For the case $(M,N)=(3,3)$, several positive periodic orbits exist; see [5]. However, it remains open whether periodic orbits exist for other cases. Still, numerical evidence suggests that positive periodic orbits exist for the cases $(M,N)=(4,4),(5,5)$; see [39].
The framework of our computer-assisted techniques (see Section 1.2) is largely model independent, and thus it paves a way to solve the problem of resource competition, for both growth rates (7)–(8). Indeed, for each fixed pair $(M,N)$ whose numerical evidence exhibits periodicity, we can employ our techniques not only to prove the existence of positive periodic orbits and thus solve the problem, but also their stability property and continuations.
1.2 General framework to desingularize transcritical bifurcations
The stable connection problem involves a continuation of positive periodic orbits through both boundary planes $Q_{1}$ and $Q_{2}$, where a transcritical bifurcation occurs. Since the validation of the continuation hinges on the local isolation of the solution, we establish a blow-up method that yields an auxiliary system of equations for which the undesired branch of periodic solutions is removed. In this section, we present this approach and explain why this strategy achieves such local isolation in the most general setting.
Consider the general system
$\partial_{t}x=f(x;\mu),$ (9)
where $f=(f_{1},\dots,f_{n})$ is analytic, $x=(x_{1},\dots,x_{n})\in\mathcal{X}\subset\mathbb{R}^{n}$, and $\mu\in U\subset\mathbb{R}$ is a parameter. We are interested in the following situation:
  1. (A1)
    There exists an invariant subset $Q\coloneqq\{x\in\mathcal{X}\,:\,G(x;\mu)=0\}$ of $\mathbb{R}^{n}$, where $G:\mathcal{X}\times U\to\mathbb{R}$ is analytic such that $\partial_{x_{j}}G(x;\mu)\neq 0$ for some $j=1,\dots,n$ and all $(x;\mu)\in\mathcal{X}\times U$, i.e., $\partial_{x_{j}}G(x;\mu)$ is invertible. Without loss of generality, we assume $j=1$.
  2. (A2)
    There exist two families $\mathcal{P}_{1}(\mu)$ and $\mathcal{P}_{2}(\mu)$ of periodic orbits parametrized by $\mu\in U$ such that $\mathcal{P}_{1}(\mu)\subset Q$ for all $\mu\in U$ and $\mathcal{P}_{2}(\mu)\cap Q=\emptyset$ for $\mu\neq\mu_{\mathrm{bif}}\in U$. Moreover, $\mathcal{P}_{1}(\mu_{\mathrm{bif}})=\mathcal{P}_{2}(\mu_{\mathrm{bif}})$, i.e., the two families meet on the invariant subset $Q$ at $\mu=\mu_{\mathrm{bif}}$.
Let $x=x(t;\mu)$ be a solution of (9). We introduce the blow-up coordinates $\zeta:U\to\mathbb{R}$ and $u:\mathbb{R}\times U\to\mathbb{R}$ such that
$G(x(t;\mu);\mu)=\zeta(\mu)u(t;\mu).$ (10)
By (A1), the implicit function theorem implies that $x_{1}=\omega(\zeta(\mu)u,x_{2},\dots,x_{n};\mu)$ for some smooth function $\omega$. Differentiating (10) with respect to $t$ yields
$\zeta(\mu)\partial_{t}u(t;\mu)=\nabla G(z(t;\mu);\mu)f(z(t;\mu);\mu),\qquad z\coloneqq(\omega(\zeta(\mu)u,x_{2},\dots,x_{n};\mu),x_{2},\dots,x_{n}),$ (11)
where $\nabla$ is the gradient with respect to the variables $x=(x_{1},\dots,x_{n})$. Also, if $t\mapsto x(t;\mu)$ represents the periodic solution to (9) associated to $\mathcal{P}_{2}$, then by (A2), we have $\mathcal{P}_{2}(\mu_{\mathrm{bif}})\subset Q$ and so $\zeta(\mu_{\mathrm{bif}})=0$. From the invariance of $Q=\{x\in\mathcal{X}\,:\,G(x;\mu)=0\}$, we get
$\nabla G(x;\mu_{\mathrm{bif}})f(\omega(0,x_{2},\dots,x_{n};\mu_{\mathrm{bif}}),x_{2},\dots,x_{n};\mu_{\mathrm{bif}})=0.$ (12)
This guarantees that $\lim_{\mu\to 0}\partial_{t}u(t;\mu)$ exists, as seen by expanding the right-hand side of (11) near $x_{1}=0$, such that
$\begin{split}\partial_{t}u(t;\mu)&=\frac{1}{\zeta(\mu)}\nabla G(z(t;\mu);\mu)f(z(t;\mu);\mu)\\ &=\sum_{\ell=1}^{\infty}c_{\ell}(x_{2}(t;\mu),\dots,x_{n}(t;\mu);\mu)\zeta(\mu)^{\ell-1}u(t;\mu)^{\ell},\end{split}$ (13)
where
$c_{\ell}(x_{2},\dots,x_{n};\mu)\coloneqq\left.\partial_{x_{1}}^{\ell}\right|_{x_{1}=0}\nabla G(z;\mu)f(z;\mu),\qquad\ell\geq 1.$ (14)
This yields the auxiliary system
$\begin{cases}\displaystyle\partial_{t}\begin{pmatrix}u\\ x_{2}\\ \vdots\\ x_{n}\end{pmatrix}=\begin{pmatrix}\frac{1}{\zeta(\mu)}\nabla G(\omega,x_{2},\dots,x_{n};\mu)f(\omega,x_{2},\dots,x_{n};\mu)\\ f_{2}(\omega,x_{2},\dots,x_{n};\mu)\\ \vdots\\ f_{n}(\omega,x_{2},\dots,x_{n};\mu)\end{pmatrix},\\ G(\omega,x_{2},\dots,x_{n};\mu)=\zeta(\mu)u,\\ u(0;\mu)=1.\end{cases}$ (15)
We have that if $(u,x_{2},\dots,x_{n})$ is a periodic solution, for some $\zeta$, to the auxiliary system (15), then $(\zeta u,x_{2},\dots,x_{n})$ is a periodic solution to the original system (9). Provided that $\zeta\neq 0$, the periodic solution does not lie in $Q$. Importantly, periodic orbits are locally isolated whenever $\zeta=0$. For completeness, we state this in the following lemma. However, we stress that verifying its hypotheses is unnecessary for our proof; it serves to justify why using the coordinates (10) and studying the auxiliary system (15) is compatible with a contraction argument (detailed in Section 2.2) to continue a branch of periodic solutions through a non-degenerate transcritical bifurcation.
Proposition 1.2 (Local isolation of periodic orbits).
Under the assumptions (A1) and (A2), the family of periodic orbits $\mathcal{P}_{2}(\mu)$, seen in the coordinates (10) and solving (15), is locally isolated near $\mu=0$.
Proof.
Let $x_{*}(t;\mu)$ denote the periodic solution associated with $\mathcal{P}_{2}(\mu)$; we need to prove local uniqueness at $\mu=\mu_{\mathrm{bif}}$. Without loss of generality, we assume $\mu_{\mathrm{bif}}=0$.
At $\mu=0$, since $Q=\{x\in\mathcal{X}\,:\,G(x;\mu)=0\}$ is an invariant subset by (A1), each $x_{j}$-equation of (15), for $j=2,\dots,n$, is decoupled from the $u$-equation. From (A2), we know that $x_{*}(t;0)$ is contained in $Q$ and solves the decoupled system (i.e., without the $u$-coordinate). Moreover, due to this decoupling, the first equation becomes a linear non-autonomous ODE with periodic coefficients
$\partial_{t}u(t;0)=\phi(t)u(t;0).$
The general solution reads $u(t;0)=Ce^{\int_{0}^{t}\phi(s)\,\mathrm{d}s}>0$, where $C=1/e$ is uniquely determined by the scaling $u(0;0)=1$. ∎
In the specific predator-prey system (5), $\mathcal{X}$ is the closure of $\mathbb{R}_{+}^{3}$, $\mu\in U=(0,\infty)$ is the carrying capacity $\kappa$, $Q=\{x\in\mathcal{X}\,:\,x_{1}=0\}$ (resp., $Q=\{x\in\mathcal{X}\,:\,x_{2}=0\}$) is the boundary plane $Q_{2}$ (resp., $Q_{1}$), $\mathcal{P}_{1}(\mu)$ is the family of boundary limit cycles, and $\mathcal{P}_{2}(\mu)$ is the continuation of positive periodic orbits. Also, the coordinates (10) reduce to $X_{1}=a_{1}u_{1}$ and $X_{2}=a_{2}u_{2}$ to handle, respectively, the two boundary planes $Q_{1}$ and $Q_{2}$. Moreover, $\omega$ is known explicitly to be $\omega=au$.